OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5). A. 23 B. –53 C. –9 D. 3

OpenStudy (anonymous):

(f ° g)(x)=(-2x - 7 )*(-4x + 3) to find (f ° g)(-5). keep x=-5 on both sides

OpenStudy (anonymous):

idk how to that

OpenStudy (anonymous):

(-2x - 7) (-4x + 3) substitute x = -5 and solve

OpenStudy (anonymous):

i keep getting 26

OpenStudy (anonymous):

(-2(-5) - 7) (-4(-5) + 3) (10 - 7) (20 + 3) = ?

OpenStudy (anonymous):

i'm getting 69 no matter what i do.

OpenStudy (anonymous):

So far, y'all have been trying to multiply, when the problem is a composition problem. I.e.,\[(f ° g)(-5)=f(g(-5))=f(23)=-53\]

OpenStudy (anonymous):

is this a composition or are you multiplying f and g

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

The circle or dot means composition.\[f \circ g(x)=f(g(x))\]